find:(f∘g)(3) where f(x)=4x+1 and g(x)=x²+1

Answer by Edwin McCravy(20056)

find:(f∘g)(3)      f(x)=4x+1 and g(x)=x²+1

(f∘g)(3) = f(g(3)) = f(3²+1) = f(9+1) = f(10) = 4(10)+1 = 40+1 = 41

Edwin